you deposit $2000 in an account that pays 4.5 % interest compounded annually. Write a model giving the account balance A after T years. In about how many years will the balance double?
Answer by stanbon(75887)

You deposit $2000 in an account that pays 4.5 % interest compounded annually. Write a model giving the account balance A after T years.
A(t) = 2000(1+(0.045))^(t)
----------------------------------
In about how many years will the balance double?
4000 = 2000(1.045)^t
2 = 1.045^t
---
t*log(1.045) = log(2)
---
t = log(2)/log(1.045)
---
t = 15.75 years
====================
Cheers,
Stan H.
===============